# Understanding Fairness and Prediction Error through Subspace Decomposition and Influence Analysis

### v1

@@ 3.1 Motivation
ensuring that $B_{z}^{\top}\boldsymbol{X}$ contains all the information in $\boldsymbol{X}$ relevant to predicting $\boldsymbol{Z}$. The column space of $B_{z}$, known as the SDR subspace of $\boldsymbol{X}$ with respect to $\boldsymbol{Z}$, see Cook and Li (2002); Adragni and Cook (2009), thus serves as a natural candidate for $\mathcal{S}_{2}$ with $\dim(\mathcal{S}_{2})=r$. Its orthogonal complement, spanned by $P=I-B_{z}B_{z}^{\top}$, defines $\mathcal{S}_{1}$ with $\dim(\mathcal{S}_{1})=p-r$. Then $h(\boldsymbol{X})=P\boldsymbol{X}$ removes all information associated with $\boldsymbol{Z}$, yielding a fair projection that satisfies $h(\boldsymbol{X})\perp\!\!\!\perp\boldsymbol{Z}$.
@@ 3.3 Utility and Fairness Trade-off
Recall that the full sufficient representation of $\boldsymbol{Y}$ is given by $U=BB^{\top}\boldsymbol{X}$, and let $\boldsymbol{\Xi}^{(m)}=P^{(m)}\boldsymbol{X}$ denote the reduced representation. The following result characterizes the prediction error of the Bayes optimal predictor based on the reduced representation.
@@
Let $\tilde{f}^{(m)}(\boldsymbol{\Xi}^{(m)})=\mathbb{E}[\boldsymbol{Y}\mid\boldsymbol{\Xi}^{(m)}]$ denote the Bayes predictor using the partially fair representation $\boldsymbol{\Xi}^{(m)}$, and let $f^{*}(U)=\mathbb{E}[\boldsymbol{Y}\mid U]$ denote the Bayes optimal predictor using the full sufficient representation. Then, the expected squared prediction error satisfies
$\mathbb{E}[(\boldsymbol{Y}-\tilde{f}^{(m)}(\boldsymbol{\Xi}^{(m)}))^{2}]=\underbrace{\mathbb{E}[\mathrm{Var}(f^{*}(U)\mid\boldsymbol{\Xi}^{(m)})]}_{\text{Approximation error}}+\underbrace{\mathbb{E}[\varepsilon_{Y}^{2}]}_{\text{Irreducible noise}}:=\Delta(m)+\sigma_{Y}^{2}.$
@@
This decomposition follows from the orthogonality principle in $L^{2}$ space, which ensures that the conditional expectation minimizes mean squared error. The term $\mathbb{E}[\mathrm{Var}(f^{*}(U)\mid\boldsymbol{\Xi}^{(m)})]$ quantifies the loss in predictive information due to reducing the representation to $\boldsymbol{\Xi}^{(m)}$. As more shared directions are included, the reduced representation becomes increasingly informative, and the error $\Delta(m)$ decreases. When $m=s$, the full central subspace for $\boldsymbol{Y}$ is recovered, yielding $\Delta(s)=0$.
@@ Appendix A Influence Function of Projections
Then, by Zhu and Fang [1996], the influence function of the directions estimated by SDR techniques can be written as, for $i=1,\ldots,s$,
@@
The influence functions for the candidate matrices $M_{Y}$ and $M_{Z}$ estimated via SDR techniques have been well studied. For details, we refer readers to Section 4 of Kim et al. [2020] and omit the derivations here.
@@ B.1 Experiment Setup
Throughout the experiments, we use the MSAVE method to estimate the candidate matrices $M_{Y}$ and $M_{Z}$, with the number of slices set to $p+1$, where $p$ denotes the dimensionality of the input $\boldsymbol{X}$. In the post SDR training procedure (as described in Algorithm 2), we apply multivariate linear regression for synthetic simulations and logistic regression for real-world datasets. The MCDP metric is used to select the optimal fair model $\mathcal{M}_{\text{fair}}^{(m^{*})}$. To estimate the rank of each candidate matrix, we adopt the ladle estimator with 30 bootstrap replications.
For baseline methods, we use the official code released by the respective authors. Specifically, for AdvDebias, FairMixup, DRAlign, and DiffMCDP, we use the implementation provided in Jin et al. [2024]. For INLP, we set the number of iterations to 100. For RLACE, we follow the training hyperparameters used in the original paper. For SUP, we adopt the same number of slices ($p+1$) as in our method, and select the dimension to remove based on 10-fold cross-validation, choosing the model that achieves the lowest MCDP.
@@ B.2 Fairness Measurement
Maximum Class-wise Discrepancy in Predictions (MCDP).
@@
When $K=2$, this definition reduces to the binary case as in Jin et al. [2024].
@@ Appendix C Proof of Theorems
$\mathbb{E}[(Y-\tilde{f}^{(m)})^{2}]=\mathbb{E}[(f^{*}(U)+\varepsilon_{Y}-\tilde{f}^{(m)})^{2}]=\mathbb{E}\left[(f^{*}(U)-\tilde{f}^{(m)})^{2}+2(f^{*}(U)-\tilde{f}^{(m)})\varepsilon_{Y}+\varepsilon_{Y}^{2}\right].$
Take expectations term-by-term. For the first term, we observe that $\tilde{f}^{(m)}=\mathbb{E}[Y\mid\boldsymbol{\Xi}^{(m)}]$, which is the projection of $Y$ onto a coarser sigma-algebra. Therefore $\tilde{f}^{(m)}=\mathbb{E}[f^{*}(U)\mid\boldsymbol{\Xi}^{(m)}]$ and
$\mathbb{E}[(f^{*}(U)-\tilde{f}^{(m)})^{2}]=\mathbb{E}[\operatorname{Var}(f^{*}(U)\mid\boldsymbol{\Xi}^{(m)})].$
For the second term, since $\mathbb{E}[\varepsilon_{Y}\mid\boldsymbol{\Xi}^{(m)}]=\mathbb{E}[\mathbb{E}[\varepsilon_{Y}\mid U]\mid\boldsymbol{\Xi}^{(m)}]=0$, we have:
$\mathbb{E}[(f^{*}(U)-\tilde{f}^{(m)})\varepsilon_{Y}]=0.$
@@
$\mathbb{E}[(Y-\tilde{f}^{(m)})^{2}]=\mathbb{E}[\operatorname{Var}(f^{*}(U)\mid\boldsymbol{\Xi}^{(m)})]+\sigma_{Y}^{2}.$
Note that $\tilde{f}^{(m)}(\boldsymbol{\Xi}^{(m)})$ is the orthogonal projection of $f^{*}(U):=\mathbb{E}[\boldsymbol{Y}\mid U]$ onto $\sigma(\boldsymbol{\Xi}^{(m)})$ in $L^{2}$. Since $\boldsymbol{\Xi}^{(m)}\subset\boldsymbol{\Xi}^{(m+1)}$, the projection becomes finer, and by the Pythagorean identity in $L^{2}$:
$\|f^{*}(U)-\tilde{f}^{(m+1)}\|_{L^{2}}^{2}\leq\|f^{*}(U)-\tilde{f}^{(m)}\|_{L^{2}}^{2}.$
Hence, the approximation error $\Delta(m):=\mathbb{E}[\operatorname{Var}(f^{*}(U)\mid\boldsymbol{\Xi}^{(m)})]$ satisfies $\Delta(m+1)\leq\Delta(m)$. ∎
@@
where, for any full rank $B$, $P_{\Sigma,B}$ is the $\Sigma$ orthogonal projector onto $\mathrm{Span}\{B\}$.

### v2

@@ 3.1 Motivation
ensuring that $B_{z}^{\top}\boldsymbol{X}$ contains all the information in $\boldsymbol{X}$ relevant to predicting $\boldsymbol{Z}$. The column space of $B_{z}$, known as the SDR subspace of $\boldsymbol{X}$ with respect to $\boldsymbol{Z}$, see Cook and Li (2002); Adragni and Cook (2009), thus serves as a natural candidate for $\mathcal{S}_{2}$ with $\dim(\mathcal{S}_{2})=r$. Its orthogonal complement, spanned by $P=I-B_{z}B_{z}^{\top}$, defines $\mathcal{S}_{1}$ with $\dim(\mathcal{S}_{1})=p-r$. Then $h(\boldsymbol{X})=P\boldsymbol{X}$ removes information associated with $\boldsymbol{Z}$, yielding a fair projection.
@@ 3.3 Utility and Fairness Trade-off
Let $\boldsymbol{\Xi}^{(m)}=P^{(m)}\boldsymbol{X}$ denote the reduced representation. The following result characterizes the prediction error of the Bayes optimal predictor based on the reduced representation.
@@
Let $\tilde{f}^{(m)}(\boldsymbol{\Xi}^{(m)})=\mathbb{E}[\boldsymbol{Y}\mid\boldsymbol{\Xi}^{(m)}]$ denote the Bayes predictor using the partially fair representation $\boldsymbol{\Xi}^{(m)}$, and let $f^{*}(\boldsymbol{X})=\mathbb{E}[\boldsymbol{Y}\mid\boldsymbol{X}]$ denote the Bayes optimal predictor using the original representation. Then, the expected squared prediction error satisfies
$\mathbb{E}[(\boldsymbol{Y}-\tilde{f}^{(m)}(\boldsymbol{\Xi}^{(m)}))^{2}]=\underbrace{\mathbb{E}[\mathrm{Var}(f^{*}(\boldsymbol{X})\mid\boldsymbol{\Xi}^{(m)})]}_{\text{Approximation error}}+\underbrace{\mathbb{E}[\varepsilon_{Y}^{2}]}_{\text{Irreducible noise}}:=\Delta(m)+\sigma_{Y}^{2}.$
@@
This decomposition follows from the orthogonality principle in $L^{2}$ space, which ensures that the conditional expectation minimizes mean squared error. The term $\mathbb{E}[\mathrm{Var}(f^{*}(\boldsymbol{X})\mid\boldsymbol{\Xi}^{(m)})]$ quantifies the loss in predictive information due to reducing the representation to $\boldsymbol{\Xi}^{(m)}$. As more shared directions are included, the reduced representation becomes increasingly informative, and the error $\Delta(m)$ decreases. When $m=s$, the full central subspace for $\boldsymbol{Y}$ is recovered, yielding $\Delta(s)=0$.
@@ Appendix A Influence Function of Projections
Then, by [52], the influence function of the directions estimated by SDR techniques can be written as, for $i=1,\ldots,s$,
@@
The influence functions for the candidate matrices $M_{Y}$ and $M_{Z}$ estimated via SDR techniques have been well studied. For details, we refer readers to Section 4 of [53] and omit the derivations here.
@@ B.1 Experiment Setup
Throughout the experiments, we use the MSAVE method to estimate the candidate matrices $M_{Y}$ and $M_{Z}$, with the number of slices set to $p+1$, where $p$ denotes the dimensionality of the input $\boldsymbol{X}$. In the post-SDR training procedure (as described in Algorithm 2), we apply multivariate linear regression for synthetic simulations and logistic regression for real-world datasets. The MCDP metric is used to select the optimal fair model $\mathcal{M}_{\text{fair}}^{(m^{*})}$. To estimate the rank of each candidate matrix, we adopt the ladle estimator with 30 bootstrap replications.
For baseline methods, we use the official code released by the respective authors. Specifically, for AdvDebias, FairMixup, DRAlign, and DiffMCDP, we use the implementation provided in [21]. For INLP, we set the number of iterations to 100. For RLACE, we follow the training hyperparameters used in the original paper. For SUP, we adopt the same number of slices ($p+1$) as in our method, and select the dimension to remove based on 10-fold cross-validation, choosing the model that achieves the lowest MCDP.
@@ B.2 Fairness Measurement
Maximal Cumulative Ratio Disparity along Predictions (MCDP).
@@
When $K=2$, this definition reduces to the binary case as in [21].
@@ Appendix C Proof of Theorems
$\mathbb{E}[(Y-\tilde{f}^{(m)})^{2}]=\mathbb{E}[(f^{*}(\boldsymbol{X})+\varepsilon_{Y}-\tilde{f}^{(m)})^{2}]=\mathbb{E}\left[(f^{*}(\boldsymbol{X})-\tilde{f}^{(m)})^{2}+2(f^{*}(\boldsymbol{X})-\tilde{f}^{(m)})\varepsilon_{Y}+\varepsilon_{Y}^{2}\right].$
Take expectations term-by-term. For the first term, we observe that $\tilde{f}^{(m)}=\mathbb{E}[Y\mid\boldsymbol{\Xi}^{(m)}]$, which is the projection of $Y$ onto a coarser sigma-algebra. Therefore $\tilde{f}^{(m)}=\mathbb{E}[f^{*}(\boldsymbol{X})\mid\boldsymbol{\Xi}^{(m)}]$ and
$\mathbb{E}[(f^{*}(\boldsymbol{X})-\tilde{f}^{(m)})^{2}]=\mathbb{E}[\operatorname{Var}(f^{*}(\boldsymbol{X})\mid\boldsymbol{\Xi}^{(m)})].$
For the second term, since $\mathbb{E}[\varepsilon_{Y}\mid\boldsymbol{\Xi}^{(m)}]=\mathbb{E}[\mathbb{E}[\varepsilon_{Y}\mid\boldsymbol{X}]\mid\boldsymbol{\Xi}^{(m)}]=0$, we have:
$\mathbb{E}[(f^{*}(\boldsymbol{X})-\tilde{f}^{(m)})\varepsilon_{Y}]=0.$
@@
$\mathbb{E}[(Y-\tilde{f}^{(m)})^{2}]=\mathbb{E}[\operatorname{Var}(f^{*}(\boldsymbol{X})\mid\boldsymbol{\Xi}^{(m)})]+\sigma_{Y}^{2}.$
Note that $\tilde{f}^{(m)}(\boldsymbol{\Xi}^{(m)})$ is the orthogonal projection of $f^{*}(\boldsymbol{X}):=\mathbb{E}[\boldsymbol{Y}\mid\boldsymbol{X}]$ onto $\sigma(\boldsymbol{\Xi}^{(m)})$ in $L^{2}$. Since $\boldsymbol{\Xi}^{(m)}\subset\boldsymbol{\Xi}^{(m+1)}$, the projection becomes finer, and by the Pythagorean identity in $L^{2}$:
$\|f^{*}(\boldsymbol{X})-\tilde{f}^{(m+1)}\|_{L^{2}}^{2}\leq\|f^{*}(\boldsymbol{X})-\tilde{f}^{(m)}\|_{L^{2}}^{2}.$
Hence, the approximation error $\Delta(m):=\mathbb{E}[\operatorname{Var}(f^{*}(\boldsymbol{X})\mid\boldsymbol{\Xi}^{(m)})]$ satisfies $\Delta(m+1)\leq\Delta(m)$. ∎
@@
where, for any full rank $B$, $P_{\Sigma,B}$ is the $\Sigma$-orthogonal projector onto $\mathrm{Span}\{B\}$.